The thickness of a printed circuit board is required to lie between the specification limits 0.150 ± 0.004 cm. A machine produces circuit boards with a thickness that is normally distributed with mean 0.151 cm and standard deviation 0.003 cm. a) What is the probability that the thickness X of a circuit board produced by this machine will fall within the specification limits? b) Now consider the mean thickness X- (X1 +X2 + ..+X25)/25 for a batch of 25 circuit boards. What is the probability that this batch mean will fall within the specification limits? (Assume that X1, X2, , X25 are independent random variables with the same distribution as X above.)
Answer #1

a) The probability that the thickness X of a circuit board produced by this machine will fall withing the specification limits is

P(0.150 ±0.004)= P(0.146 <X <0.154)-Pl -40-μ (0.146-0.151 Ζ<0.154-0.151 0.003 0.003 = P(-1.6667 <Z <1) = P(Z <1)-P(Z <-1.667) = 0.84134-0.04776=0.79358 from normal distribution tables

b) The probability that the thickness Xbar of a circuit board produced by this machine will fall withing the specification limits is

P(0.150±0.004)- p(0. 146-0. 151ノワ. O154-0. 151 0.003/250003/25 = P(-8.3333 < Z < 5) = P(Z < 5)-P(Z <-8.333) 1-0 = 1、from normal distribution tables